Cooperton Mining just announced it will cut its dividend from

$4.17

to

$ 2.36

per share and use the extra funds to expand. Prior to the​ announcement, Cooperton's dividends were expected to grow at a

3.2 %

​rate, and its share price was

$ 50.68

With the planned​ expansion, Cooperton's dividends are expected to grow at a

4.5 %

rate. What share price would you expect after the​ announcement? (Assume that the new expansion does not change​ Cooperton's risk.) Is the expansion a good​ investment?
Answer #1

Rate of return can be computed using GGM for price of stock as:

P0 = D1/ (r -g)

D1 = Expected dividend = $ 4.17 x 1.032 = $ 4.30344

r = Rate of return

g = Growth rate in dividend = 0.032

$ 50.68 = $ 4.30344/ (r -0.032)

r -0.032 = $ 4.30344/$ 50.68

             = 0.08491397

r = 0.08491397 + 0.032 = 0.11691397

Expected price of share after dividend cut,

P = $ 2.36 x (1+0.045)/ (0.11691397– 0.045)

= $ 2.36 x (1.045)/ (0.11691397– 0.045)

= $ 2.4662/ 0.07191397

= $ 34.29375405 or $ 34.29

Expected price of share after the announcement is $ 34.29

Expansion is not a good investment as it decreased the share price.